# Alternating Bias Assisted Annealing of Amorphous Oxide Tunnel Junctions

### 15 Jan 2024 v1

@@ Abstract
We demonstrate a transformational technique for controllably tuning the electrical properties of fabricated thermally oxidized amorphous aluminum-oxide tunnel junctions. Using conventional test equipment to apply an alternating bias to a heated tunnel barrier, giant increases in the room temperature resistance, greater than 70%, can be achieved. The rate of resistance change is shown to be strongly temperature-dependent, and is independent of junction size in the sub-micron regime. In order to measure their tunneling properties at mK temperatures, we characterized transmon qubit junctions treated with this alternating-bias assisted annealing (ABAA) technique. The measured frequencies follow the Ambegaokar-Baratoff relation between the shifted resistance and critical current. Further, these studies show a reduction of junction-contributed loss on the order of , along with a significant reduction in resonant- and off-resonant-two level system defects when compared to untreated samples. Imaging with high-resolution TEM shows that the barrier is still predominantly amorphous with a more uniform distribution of aluminum coordination across the barrier relative to untreated junctions. This new approach is expected to be widely applicable to a broad range of devices that rely on amorphous aluminum oxide, as well as the many other metal-insulator-metal structures used in modern electronics.
@@ I Introduction
The preparation of amorphous thin film materials is surprisingly simple and devices based on them are ubiquitous in electronics, but their morphology is surprisingly complex and challenging to control. Among other similar materials, amorphous aluminum oxide tunnel (a-AlO) junctions stand out due to their importance in both cryogenic and room temperature devices. Applications range from those that utilize the Josephson effect Josephson (1962), such as SQUID magnetometers Clarke (2004), superconducting qubits Devoret (1997); Nakamura et al. (1999), amplifiers Castellanos-Beltran et al. (2009); Macklin et al. (2015), RSFQ logic Likharev and Semenov (1991), to room temperature magnetic tunnel junctions and memory elements Zhu and Park (2006); Rippard et al. (2002); O’Dwyer (2004).
@@
Unfortunately, while it is essential that , and hence , is well defined, a large spread of in nominally identical junctions is typically observed, on the order of 5-10% at the wafer level Muthusubramanian et al. (2023). This is due to both intrinsic properties and extrinsic effects. Intrinsic issues include the exponential dependence of the tunneling on the thickness Zeng et al. (2016); grain structure; and interface roughness. Extrinsic effects arise from the effects of lithography uncertainties Bilmes et al. (2021), processing, electrostatic sensitivity, and circuit design.
@@
There have been many efforts to improve the properties of the a-AlO material, focused primarily on understanding and optimizing the junction formation and post-processing, i.e. pressure, time, Mallison (1995), interface preparation, etches, simulations, etc., see Ref. Kim et al. (2020); Cyster et al. (2021) and citations therein. These efforts have met with limited success, and it is typically assumed that the properties of the junctions, as formed, are relatively frozen-in and difficult to adjust. For example, annealing the junctions to high temperatures can slightly change the resistance, globally heating many junctions has limited efficacy due to nanoscopic morphology differences of nominally identical junctions at the domain and atomic level. Since the resistance tends to increase on continued annealing, novel approaches to sequentially targeting the junctions have been useful in mitigating these spreads by targeting a slightly lower resistance than desired and then heating the individual junctions locally to bring them into tolerance Hertzberg et al. (2021); Zhang et al. (2022b); Kim et al. (2022). However, these procedures involve the addition of optical elements to the test and measurement apparatus and appear to be limited to adjustments of the resistances up to about 10-12%, i.e. on the order of the junction spread. This limits the number of junction arrays that can be successfully adjusted, based on the initial distribution of the junctions.
On the other hand, it has been shown that oxidation can be enhanced by an external voltageKonkin and Adler (2008), for example, that generated by e-beam bombardment at the surface Ebinger and Yates (1998). When comparing with other physical material systems, e.g. magnetic Zhang (2001) and piezo-electricYamamoto et al. (2014), where annealing while applying an alternating field polarity can have significant positive effects, this approach has not yet been investigated in connection to a-AlO tunnel junctions and is illustrated in Fig 1(b).
@@ II Experimental setup and results on junctions
Further studies including the temperature dependence of this effect are shown in Fig. 3. We find the rate increases significantly as the temperature is increased from 25 up to 80C. The increase is on the order of 2 every 10C, consistent with that expected from an Arrhenius law. On the other hand, the size dependence (not shown) is flat over more than an order of magnitude, indicating that the mechanism is not due to the perimeter or other side effects. This points to local effects such as reordering and increased coordination of the atoms in the barrier, leading to changes in the effective barrier height as atoms reach a more stable configuration or interfacial modifications such as smoothing or thickening of the barrier. Furthermore, since the data was taken at constant voltage, we infer that it is not a current-driven diffusion effect. This has more practical implications from the perspective of SQUID-based devices, such as sensors and tunable qubits, because these devices are typically made using two parallel circuits with junctions that may or may not be the same size. Finally, we note that the resistance jumps are also evident in the 25, 40, and 60C data.
@@ III Effect of ABAA on transmon qubits
We probe the effect of ABAA on the tunneling properties of the barrier by characterizing transmon qubits taken from the same wafer with and without ABAA. We measure a variety of properties associated with the junctions: whether follows the behavior expected by the Ambegaokar-Baratoff relationship, the impact of the barrier on qubit loss, and the presence of strongly coupled TLS’s. We measured both tunable and fixed-frequency qubits Manenti et al. (2021). For this test, we select 5 equivalent chips. Two chips containing 16 qubits each were treated with ABAA. Of these, 24 qubits had AB voltage pulses applied (3 qubits failed to a short circuit, 8 tunable qubits lost tunability), while 8 qubits on the same chip were not subjected to AB voltage pulsing, and one of these was not measurable. These devices were characterized contemporaneously with another reference set of one chip heated to 80C for 1 hour to mimic ABAA processing (16 qubits, one of which was not measurable), and two chips were not post-processed with annealing (32 qubits, 2 were not measurable).
@@
The Ambegaokar-Baratoff parameter set for this ensemble was first determined using the untreated devices. Then the measured frequencies of the treated qubits were calculated for various resistance-trimming amounts. For the non-tunable qubits, this is straightforward, while for the tunable qubits, it is necessary to find the maximum frequency, i.e. zero flux-bias, point. This process is illustrated in Fig 4(a), showing the spectrum of a tunable qubit. The spectrum not only allows one to find the maximum frequency, but it also probes for the existence of any strongly coupled TLS in this frequency range. In fact, for this qubit and all of the other tunable devices in the ABAA ensemble, for a total tunability range of GHz in 9 tunable qubits, we found no evidence of strongly coupled TLSs. Measurements on reference qubits that did not receive any additional processing, and those that were heated to C but not ABAA processed have TLS/GHz and TLS/GHz, respectively. Furthermore, the low-frequency stability of one of the devices was monitored over many hours and there was no evidence of coupling to TLFs Schlör et al. (2019).
@@
In Figure 5 we show coherence metrics for all qubits measured in the study. Here, each data point corresponds to the median of many measurements on one particular qubit. We observe that the ABAA processed devices have a higher median than both the unprocessed and the devices that were only heated to C, although the and values between sample sets are not distinguishable within the spread of the values, as shown in Table 1.
Qubit Coherence | (s) | (s) | (s) |
ABAA treated | |||
80 C annealed | |||
Unprocessed |
@@
When analyzing processes that may reduce loss in qubits, it is useful to compare these metrics in the form of a frequency-normalized loss tangent (); the ABAA tuned sample set has median loss tangent of , while both the C annealed and unprocessed sets have median loss tangents of and , respectively. We found that the coherence was typically improved after the ABAA process, indicating that the resistance trimming may have reduced the median loss tangent in the junction by about . Furthermore, the effects of the resistance trimming on the qubit frequencies are shown in Fig. 4(b), where we see that the increase of the resistance results in a reduced measured frequency, shown in Fig. 4(c) as orange stars. The expected frequencies (blue circles), compare well with the measured values. We also find the Ambegaokar-Baratoff relationship is consistent between junctions with ABAA applied vs. unprocessed with values of pH/ and pH/ respectively.
Finally, TEM imaging was employed to investigate morphological changes in the junctions due to the ABAA process. Secondary electron images for the JJ surface were acquired using a scanning electron microscope (SEM) (Helios, Thermo Fisher Scientific Ltd.) at 2 kV. Transmission electron microscope (TEM) samples were prepared by a focused ion beam instrument with a gas injection system (Helios, Thermo Fisher Scientific Ltd.). The TEM samples were thinned to electron beam transparency by a Ga ion beam from 30 to 2 kV. The TEM samples were investigated by an aberration-corrected TEM (Titan Cube, Thermo Fisher Scientific Ltd.) at 200 kV. A high-angle annular dark-field (HAADF) detector was used for dark-field imaging in scanning TEM (STEM) mode with a convergent semi-angle and a collection semi-angle of 18 mrad and 74-200 mrad, respectively. Energy-dispersive X-ray spectroscopy (EDS) and electron energy-loss spectroscopy (EELS) studies were carried out with probe currents of 250 nA and 50 nA, respectively. Dual EELS was performed to acquire all-electron energy-loss (ELL) spectra. The plural scattering effect of all the raw ELL spectra was removed via Fourier-ratio deconvolution.
@@
TEM imaging and EELS analysis of the ABAA-treated sample indicate that the increase in junction resistivity may be attributed to a more uniform barrier structure with reduced point defect density. An initial survey of the barrier thickness across each device showed no apparent difference in thickness between processed and unprocessed junctions, and the barrier remains amorphous after processing. A low-mag cross-sectional TEM/STEM image and corresponding EDS elemental distribution map are shown in Fig. 6(a-c). An amorphous a-AlO barrier is observed in the ABAA-treated sample (Fig. 6(d)). HAADF-STEM image shows a barrier thickness of approximately 2 nm (Fig. 6(e)).
@@ Acknowledgements.
The structural and chemical characterization and analysis is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract number DE-AC02-07CH11359. We acknowledge the support of Anna Grassellino and Akshay Murthy. All resistance measurements were conducted on FormFactor automated probe stations with the help of Anwar Aslam, Connor Smith, and Brandon Boiko. We thank Hilal Cansizoglu and the Rigetti fab team members for junction process development and samples, as well as Rory Cochrane for cryogenic measurements of transmon devices at mK temperatures. We thank Greg Stiehl for his critical eye in preparation of this manuscript.References

### 27 Feb 2024 v2

@@ Abstract
We demonstrate a transformational technique for controllably tuning the electrical properties of individual aluminum-oxide tunnel junctions. Using conventional test equipment to apply an alternating bias to a heated tunnel barrier, giant increases in the room temperature resistance, greater than 70%, can be achieved. The rate of resistance change is shown to be strongly temperature-dependent, and is weakly dependent on junction size in the sub-micron regime. In order to measure their tunneling properties at mK temperatures, we characterized transmon qubit junctions treated with this alternating-bias assisted annealing (ABAA) technique. The measured frequencies follow the Ambegaokar-Baratoff relation between the shifted resistance and critical current. Further, these studies show a reduction of junction-contributed loss on the order of , and indicate that there is a reduction in resonant-and off-resonant-two level system defects when compared to untreated samples. Imaging with high-resolution TEM shows that the barrier is still predominantly amorphous with a more uniform distribution of aluminum coordination across the barrier relative to untreated junctions. This new approach is expected to be widely applicable to a broad range of devices that rely on amorphous aluminum oxide, and is likely applicable to many other metal-insulator-metal devices.
@@ I Introduction
The preparation of amorphous thin film materials is surprisingly simple. While devices incorporating these materials are ubiquitous, their morphology is surprisingly complex and challenging to control. Among other similar materials, amorphous aluminum oxide tunnel (a-AlO) junctions stand out due to their importance in both cryogenic and room temperature devices. These applications range from those that utilize the Josephson effect Josephson (1962), such as SQUID magnetometers Clarke (2004), superconducting qubits Devoret (1997); Nakamura et al. (1999), amplifiers Castellanos-Beltran et al. (2009); Macklin et al. (2015), RSFQ logic Likharev and Semenov (1991), to room temperature magnetic tunnel junctions Shajil Nair et al. (2023) and memory elements Zhu and Park (2006); Rippard et al. (2002); O’Dwyer (2004).
@@
Unfortunately, while it is essential that , and hence , is well defined, a large spread of in nominally identical junctions is typically observed, on the order of 1-10% at the wafer level Muthusubramanian et al. (2023). This is due to both intrinsic properties and extrinsic effects. Intrinsic issues include the exponential dependence of the tunneling on the thickness Zeng et al. (2016); grain structure; and interface roughness. Extrinsic effects arise from the effects of lithography uncertainties Bilmes et al. (2021), processing, electrostatic sensitivity, and circuit design.
@@
There have been many efforts to improve the properties of the a-AlO material, focused primarily on understanding and optimizing the junction formation and post-processing, i.e. pressure, time, Mallison (1995), interface preparation, etches, simulations, etc., see Ref. Kim et al. (2020); Cyster et al. (2021) and citations therein. These efforts have met with limited success, and it is typically assumed that the properties of the junctions, as formed, are relatively frozen-in and difficult to adjust. For example, annealing the junctions to high temperatures can significantly change significant the resistanceAnnealingNbAlAlOxNbVettoliere2020, globally heating many junctions has limited efficacy due to nano-scopic morphology differences of nominally identical junctions at the domain and atomic level. Since the resistance tends to increase on continued annealing, novel approaches to sequentially targeting the junctions have been useful in mitigating these spreads by targeting a slightly lower resistance than desired and then locally heating the individual junctions to bring them into tolerance Granata et al. (2007); Hertzberg et al. (2021); Zhang et al. (2022b); Kim et al. (2022). However, these procedures involve the addition of optical elements to the test and measurement apparatus and appear to be limited to adjustments of the resistances up to about 10-12% in Al/a-AlO/Al junctions, i.e. on the order of the junction spread. This limits the number of junction arrays that can be successfully adjusted, based on the initial distribution of the junctions.
On the other hand, it has been shown that oxidation can be enhanced by an external voltage Konkin and Adler (2008), for example, that generated by e-beam bombardment at the surface Ebinger and Yates (1998). When comparing with other physical material systems, e.g. magnetic Zhang (2001) and piezo-electric Yamamoto et al. (2014), where annealing while applying an alternating field polarity can have significant positive effects Fontanini et al. (2022); Shajil Nair et al. (2023), this approach has not yet been investigated in connection to a-AlO only tunnel junctions and is illustrated in Fig 1(b).
@@ II Experimental setup and results on junctions
Further studies including the temperature dependence of this effect are shown in Fig. 3. We find the rate increases significantly as the temperature is increased from 25 up to 80C. The increase is on the order of 2 every 10C, consistent with that expected from an Arrhenius law. On the other hand, the size dependence (shown in the supplement) is flat over more than an order of magnitude, indicating that the mechanism is not due to the perimeter or other side effects. This points to local effects such as reordering and increased coordination of the atoms in the barrier, leading to changes in the effective barrier height as atoms reach a more stable configuration or interfacial modifications such as smoothing or thickening of the barrier. Furthermore, since the data was taken at constant voltage, we infer that it is not a current-driven diffusion effect. This has more practical implications from the perspective of SQUID-based devices, such as sensors and tunable qubits, because these devices are typically made using two parallel circuits with junctions that may or may not be the same size. Finally, we note that the resistance jumps are also evident in the 25, 40, and 60C data.
@@ III Effect of ABAA on transmon qubits
We probe the effect of ABAA on the tunneling properties of the barrier by characterizing transmon qubits taken with and without ABAA. We measure a variety of properties associated with the junctions: whether follows the behavior expected by the Ambegaokar-Baratoff relationship, the impact of the barrier on qubit loss, and the presence of strongly coupled TLS’s. We measured both tunable and fixed-frequency qubits Manenti et al. (2021). For the coherence and TLS tests, we selected 5 equivalent chips from the same wafer, each containing nominally 16 qubits; 14 flux-tunable and two fixed qubits. Two chips were control devices with no ABAA process (unprocessed). Of the three remaining chips, all were processed to 80C, but qubits on one and a half chips only (24 in total) were processed with ABAA. Among the 24 ABAA processed qubits, 3 failed to a short circuit and 6 others lost tunability (we plan to understand failure mode as part of future work). Among qubits in the control group or those that were processed through 80C but not ABAA, we had a couple failed qubits each at cryogenic temperatures which showed expected room-temperature resistance measurements, possibly due to handling related issues.
@@
The Ambegaokar-Baratoff parameter set for this ensemble was first determined using the untreated devices. Then the measured frequencies of the treated qubits were calculated for various resistance-trimming amounts. For the non-tunable qubits, this is straightforward, while for the tunable qubits, it is necessary to find the maximum frequency, i.e., zero flux-bias point. This process is illustrated in Fig 4(a), showing the spectrum of a tunable qubit. It can also be seen in Fig 4(b) that the predicted and measured frequencies for ABAA treated qubits match well, as discussed more in depth later.
The qubit spectroscopy also probes for the existence of avoided level crossings in this frequency range. Such crossings are signatures of strongly coupled TLS defects within the tunnel barrier. Bilmes et al. (2022); Lisenfeld et al. (2019) For reference, measurements on unprocessed qubits show about TLS/GHz, and on those that were heated to C have TLS/GHz.
Compared to these measurements, for this qubit and all of the other tunable devices in the ABAA-treated ensemble we found no evidence of avoided level crossings. Over the total frequency range of GHz in 9 tunable qubits, we would have expected at least a few avoided-level TLS signatures based on the above reference data. Furthermore, the low-frequency stability of one of the devices was monitored over many hours and there was no evidence of coupling to TLFs Schlör et al. (2019), as illustrated in the supplement. We note here that smaller splittings, typically detected with swap spectroscopy Lisenfeld et al. (2019) and likely due to surface defects, would not be affected by the ABAA. While this small data set is not conclusive evidence that such defects are suppressed by the ABAA process, it does point to the possibility that ABAA may improve qubit coherence and frequency stability. In order to investigate this possibility further, we studied the loss and time domain coherence in the ABAA qubit ensemble.
@@
In Figure 5 and tabulated in the supplement, we show coherence metrics for all qubits measured in the study. Here, each data point corresponds to the median of at least 60 measurements on one particular qubit. We observe that the ABAA processed devices have a higher median than both the unprocessed and the devices that were only heated to C, although the and values between sample sets are not distinguishable within the spread of the values.
@@
When analyzing processes that may reduce loss in qubits, it is useful to compare these metrics in the form of a qubit decay rate normalized to frequency (); the ABAA tuned sample set has median loss tangent of , while both the C annealed and unprocessed sets have median loss tangents of . We found that the coherence was typically improved after the ABAA process, indicating that the resistance trimming may have reduced the median loss tangent in the junction by about . Furthermore, the effects of the resistance trimming on the qubit frequencies are shown in Fig. 4(b), where we see that the increase of the resistance results in a reduced measured frequency, shown in Fig. 4(c) as orange stars. The expected frequencies (blue circles), compare well with the measured values. We also find the Ambegaokar-Baratoff relationship is consistent between junctions with ABAA applied vs. unprocessed with values of pH/ and pH/ respectively.
@@
Finally, TEM imaging and EELS analysis were employed to investigate morphological changes in the junctions due to the ABAA process. These studies indicate that the increase in junction resistivity may be attributed to a more uniform barrier structure with reduced point defect density. A low-mag cross-sectional TEM/STEM image and corresponding EDS elemental distribution map are shown in Fig. 6(a-c) showing that a-AlO barrier is still amorphous in the ABAA-treated sample (Fig. 6(d)), and the HAADF-STEM image shows a barrier thickness of approximately 2 nm (Fig. 6(e)).
@@ Acknowledgements.
The structural and chemical characterization and analysis is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract number DE-AC02-07CH11359. We acknowledge the support of Anna Grassellino and Akshay Murthy. All resistance measurements were conducted on FormFactor automated probe stations with the help of Anwar Aslam, Connor Smith, and Brandon Boiko. We thank Hilal Cansizoglu and the Rigetti fab team members for junction process development and samples, as well as Rory Cochrane and Cassidy Berk for assistance with cryogenic measurements and TLS analysis. We thank Greg Stiehl for his critical eye in preparation of this manuscript.References
@@
- Shajil Nair et al. (2023) Keerthana Shajil Nair, Marco Holzer, Catherine Dubourdieu, and Veeresh Deshpande, “Cycling Waveform Dependent Wake-Up and ON/OFF Ratio in Al2O3/Hf0.5Zr0.5O2 Ferroelectric Tunnel Junction Devices,” ACS Applied Electronic Materials 5, 1478–1488 (2023).
@@
- Pauling (1931) Linus. Pauling, “The nature of the chemical bond. application of results obtained from the quantum mechanics and from a theory of paramagnetic susceptibility to the structure of molecules,” Journal of the American Chemical Society 53, 1367–1400 (1931).
@@
- Granata et al. (2007) C. Granata, A. Vettoliere, L. Petti, M. Rippa, B. Ruggiero, P. Mormile, and M. Russo, “Localized laser trimming of critical current in niobium based josephson devices,” Applied Physics Letters 90 (2007), 10.1063/1.2746060.
@@
- Fontanini et al. (2022) Riccardo Fontanini, Justine Barbot, Mattia Segatto, Suzanne Lancaster, Q. Duong, Francesco Driussi, L. Grenouillet, L. Triozon, Jean Coignus, Thomas Mikolajick, S. Slesazeck, and D. Esseni, “Interplay between charge trapping and polarization switching in beol-compatible bilayer ferroelectric tunnel junctions,” IEEE Journal of the Electron Devices Society 10, 1–1 (2022).
@@
- Bilmes et al. (2022) Alexander Bilmes, Serhii Volosheniuk, Alexey V. Ustinov, and Jürgen Lisenfeld, “Probing defect densities at the edges and inside josephson junctions of superconducting qubits,” npj Quantum Information 8 (2022), 10.1038/s41534-022-00532-4.
- Lisenfeld et al. (2019) Jürgen Lisenfeld, Alexander Bilmes, Anthony Megrant, Rami Barends, Julian Kelly, Paul Klimov, Georg Weiss, John M. Martinis, and Alexey V. Ustinov, “Electric field spectroscopy of material defects in transmon qubits,” npj Quantum Information 5 (2019), 10.1038/s41534-019-0224-1.